8 water waves hit a beach in 2 seconds. The distance between two peaks was 1.3m. Calculate the speed of the wave
1
Expert's answer
2018-09-10T12:41:08-0400
T=2/8=0.25 s.
The speed of the wave:
v=λ/T=1.3/0.25=5.2 m/s.